Two stars with masses 2.75 x 1030 kg and 5.92 x 1030 kg form a binary system. If the two stars have an average separation of 5.96
Answer #1

Here ,

gravitational force between the stars = Gm1 *m2/d^2

Fg = 6.67 *10^-11 * 2.75 *10^30 * 5.92 *10^30/(5.96 *10^11)^2

Fg = 3.06 *10^27 N

the magnitude of the gravitational force in Newtons between the stars. is 3.06 *10^27 N